Circular No. 8473

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005Q, 2005R, 2005S, 2005T, 2005U Several apparent new supernovae have been reported, all on unfiltered

CCD images except for 2005U, which was found by S. Mattila (Stockholm Observatory), R. Greimel (Isaac Newton Group of Telescopes), and C. Gerardy and W. P. S. Meikle (Imperial College, London) on K_{s}-band images obtained as part of the ongoing ‘Nuclear Supernova Search’ campaign with the William Herschel Telescope (+ LIRIS). SN 2005Q was found by L. A. G. Monard (cf. IAUC 8454), and SN 2005R was found by T. Boles (cf. IAUC 8469), while 2005S and 2005T were found by H. Pugh, J. Graham, and W. Li (LOSS/KAIT; cf. IAUC 8472).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005Q Jan. 28.803 1^{h}30^{m}03.^{s}51 −42^{o}40′48′′.4 17.2 22′′ W, 20′′ N
2005R Jan. 13.971 11 14 56.81 +33 49 35.7 17.4 0′′.9 W, 3′′.2 N
2005S Jan. 22.58 14 08 28.84 + 7 03 27.2 19.1 5′′.6 W, 0′′.9 S
2005T Jan. 24.54 14 26 17.24 +26 14 47.8 18.4 7′′.2 W, 0′′.9 N
2005U Jan. 30.30 11 28 33.22 +58 33 42.5 16.2 3′′.7 W, 4′′.9 S

Additional approximate unfiltered CCD magnitudes, by the respective discoverers unless noted otherwise: SN 2005Q in ESO 244-G31, 2004 Dec. 30.814 UT, [18.7; Jan. 29.757, 17.1 ± 0.2. SN 2005R in UGC 6274, 2004 Apr. 22, [19.5; Nov. 13, [19.5; 2005 Jan. 16.052, 17.4. SN 2005S in UGC 9037, 2004 June 24.21, [20.0; 2005 Jan. 24.59, 19.2. SN 2005T in IC 4423, 2004 June 27.21, [19.5; 2005 Jan. 31.53, 18.3. SN 2005U in Arp 299, 2004 Dec. 25.41, [18.8 (Li, KAIT); Jan. 23.40, 18.2 (Li, KAIT).

Monard adds that nothing is visible at the location of SN 2005Q on a Digitized Sky Survey image (limiting red mag 20.5). Boles remarks that nothing is visible at the location of SN 2005R on Digitized Sky Survey plates from 1992 (limiting red mag 21.0), 1990 (limiting blue mag 20.5), or 1983 (limiting visual mag 18.5); he adds that comparison with the Konkoly Observatory plate of the foreground variable designated 1985J (cf. IAUC 4059) shows that 2005R is 15′′ southwest of 1985J. Li provides the following position end figures for SN 2005U: 33^{s}.13, 41′′.3. Mattila et al. note that their offset above for SN 2005U is with respect to the K_{s}-band nucleus A of the luminous infrared galaxy Arp 299 (the tabulated magnitude above for 2005U is also in the K_{s} band); they estimate preliminary J − K = +0.4 ± 0.5 for 2005U, adding that the color indicates modest extinction.

2005 January 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green